Grade 11Physical Chemistry

a 4.5 g sample of pyrolusitewhen heated with conc.hcl gave chlorine.the chlorine thenpassed through ki solution liberated 6.35gof iodine.the percentage purity of pyrolusite sample is???

prachi
9 Years agoGrade 11

1 Answer

Vikas TU
9 Years ago
% purity is the percentage of the material which is the actually desired chemical in a sample of it.
    MASS of USEFUL PRODUCT
PERCENT PURITY   =  100  x    ------------------------------------------------------
    in TOTAL MASS of SAMPLE

Here total sample is 4.5 + Mass of  Hcl =
4.5 + 1+ 17 = 22.5
And iodine liberated in 6.35 . so now substitiuting the values = 100 x "6.35 /22.5" = 28.2 gm%
% purity is the percentage of the material which is the actually desired chemical in a sample of it.
    MASS of USEFUL PRODUCT
PERCENT PURITY   =  100  x    ------------------------------------------------------
    in TOTAL MASS of SAMPLE

Here total sample is 4.5 + Mass of  Hcl =
4.5 + 1+ 17 = 22.5
And iodine liberated in 6.35 . so now substitiuting the values = 100 x "6.35 /22.5" = 28.2 gm%